# The Superstatistics of confined ants: From serenity to panic

### v2

@@ Abstract
Nonequilibrium systems with strong parameter fluctuations are difficult to describe with ordinary Statistical Mechanics. Superstatistics –that can be seen as statistics of an statistics– has emerged as a tool able to describe these complex systems. It has been successfully applied, for example, to hydrodynamic turbulence, internal convection, and even DNA architecture, but not to macroscopic biological systems. Here we present fully controlled laboratory experiments in confined ants both in normal and in “panic” conditions. Using a data-driven superstatistical model, we reproduce key signatures of their motion: from non-Gaussian velocity distributions driven by a Log-normal diffusion coefficient, to clustering away from the source of danger information.
@@ I Significance Statement
We introduce Superstatistics as a study tool in the animal world, and report its Log-normal variant in a living system. Our study suggests that “panicked” ants, when exposed to a threat, can transmit information to their nestmates through a barrier they cannot physically cross. We explain both normal and panic dynamics, identifying non-Gaussian velocity densities, time scale separation, and the statistics of the diffusion coefficient. We also reveal distinct behavioral regimes in the ants’ collective panic response and how it relates with the individual ant memory and cluster formation. Furthermore, our findings indicate that optical signals or simple antennation are not significant mechanisms for panic transmission. We believe that these discoveries provide a foundation to uncover the biological origin of Log-normal type diffusion in confined environments.
@@ II Introduction
Nature displays a diversity of collective biological behaviors ranging from the microscopic to the macroscopic scales, and from high confined environments to the open space: the movement of cells during wound heal Vedel et al. (2013); Angelini et al. (2011); Zhang et al. (2010) and the flocking of birds Cavagna et al. (2010); Ballerini et al. (2008); Pearce et al. (2014) illustrate quite well those scenarios.
Here, we focus on the collective motion of confined macroscopic individuals, which is a common scenario in eusocial organisms, such as ants, who often live in high density societies. Leaf-cutter ants like Atta Insularis build complex nests made of networks of tunnels and chambers within which they live, move, take care of eggs, and even harvest fungus. While there is an extensive body of work that qualitatively describes the activity of ants in both open and confined spaces, its detailed quantitative understanding is still challenging Hölldobler and Wilson (1990); Gordon (2010); Sumpter (2006).
In terms of quantitative modelling, one may start invoking the Langevin equation, which captures the finding that animals tend to maintain their direction of motion during some characteristic (persistence) time Méndez et al. (2016); Viswanathan et al. (2011); Gurarie et al. (2017); Santana-Filho et al. (2020). Namely,
@@
where $v_{i}$ denotes the $i$-th component of velocity; $\gamma$, a positive parameter that encloses directional persistence; and $\sigma$, the intensity of a Gaussian white noise, $\xi_{i}$. A more complete overview of the statistical properties of Equation (1) was made in 1931 and it is known as Ornstein-Uhlenbeck process Uhlenbeck and Ornstein (1930), nowadays a reference in ecological models for persistent movement Méndez et al. (2016).
In addition to this, a superposition of different dynamics on different time scales often describes properties of driven nonequilibrium systems Beck (2001), e.g., the fast jiggly dynamics of a Brownian particle subject to slow environmental fluctuations of a given diffusivity Wang et al. (2009); Chechkin et al. (2017). This is known as Superstatistics (which stands for superposition of statistics), and one of its key merits is the ability to explain the emergence of non-Gaussian densities Beck and Cohen (2003); Metzler (2020).
The theory is connected to Langevin systems and, within the framework of Statistical Mechanics, can be interpreted by considering a system that is composed of small regions or cells that are temporarily in local equilibrium, showing spatio-temporal fluctuations of some intensive quantity, such as the inverse temperature $\beta=2\gamma/\sigma^{2}$ in Equation (1). It is assumed that the system reaches local equilibrium on a timescale $\tau_{1}$ much smaller than the time $\tau_{2}$ needed for $\beta$ to vary among cells. Consequently, within each region, $\beta$ could be considered approximately constant. Then, during time intervals of the order of $\tau_{2}$ and inside each cell of a system following the dynamics of Equation (1), local Gaussian velocity stationary distributions of the form:
@@
should be computed. Here, $f(\beta)$ is the probability density function of $\beta$, and a relation such as Equation (3) explains the emergence of many non-Gaussian behaviors reported in the literature Queirós (2008); Rizzo et al. (2004); Porporato et al. (2006); Briggs and Beck (2007); Chen and Beck (2008).
By performing careful experiments, here we use the behavior of confined ants in normal and panic conditions as a case study for superstatistical analysis. We quantify the motion of ants that are confined into a squared, quasi-2D cell. A number of interesting features arise: from non-Gaussian ant velocity distribution heavily influenced by speed-dependent noise intensity, to the existence of a fluctuating diffusivity, characteristic of non-Gaussian diffusion emergence in single-particle tracking experiments performed in soft and biological matter Wang et al. (2012); Chen and Beck (2008). Then, we expose the confined ants to the effect of danger information: ants in a neighbouring cell that are chemically excited transmit “panic information” to the individuals under study through a specially-designed barrier that cannot be physically crossed. A collective output is built within minutes: in spite of their short individual memory Gordon (2010), the ants form a cluster whose center shifts away from the barrier through which the danger information has emerged. In other words, the individual short-term memory has become a long-term collective memory when danger information is turned on. We propose a superstatistical model able to quantitatively reproduce all the experimentally observed features, both before and after panic. To the best of our knowledge, this is the first documented occurrence of Superstatistics in the animal world and Log-normal Superstatistics in any living system.
@@ III Experimental details
A system consisting in two 3D-printed cells, each of $6\times 6\times 1\,\mathrm{cm^{3}}$ and separated by a barrier, was designed to investigate the possible transmission of danger information by ants from one cell to another (see Figure 1). Ants are placed in both cells and filmed during $5$ minutes. Then, an insect repellent is added at the center of the left cell and another $10$ minutes of film is taken. The left and right cells will be referred to as DANGER and SAFE, respectively. In a CONTROL experiment the same procedure is followed, except no ants are introduced into the DANGER arena. We hypothesize that, if ants in the SAFE arena move away from the barrier after the addition of the repellent a significantly larger distance than they do during a CONTROL experiment, danger information has been transmitted by the “panicked” ants in the DANGER cell to the ants in the SAFE cell beyond any “passive” diffusion of the repellent through the barrier.
@@
In an attempt to isolate the possible mechanisms of danger information transmission, various types of barriers were employed (lower panel of Figure 1A). The glass barrier consists of a transparent glass wall aimed at filtering optical signals as the primary mechanism for the transfer of panic information. The thick barrier is made in such a way that ants on both sides can make contact with their extremities (e.g., their respective antennas) but cannot antennate the bodies of individuals in the opposite arena. In contrast, the thin barrier allows such contacts (see Figure 1C). We term our protocols GLASS, THICK and THIN, corresponding to each type of barrier.
@@ IV Collective motion: Tracking the Center of Mass
Using image processing techniques, we tracked the center of mass (CM) in both cells for each protocol. Data was processed using yupi, a Python library to handle (tracking, transform, generate, store, analyze, visualize) trajectory data Reyes et al. (2023). As illustrated in Figure 2A-C for typical experiments, the motion of the CM at the DANGER cell is always near its geometrical center, independently from the protocol used. A similar situation holds for the SAFE cell, with the conspicuous exception of THIN, where the CM tends to move away from the barrier after the repellent is added into the DANGER arena. Figure 2E-H shows the time evolution of the $x$-coordinate of the CM at the SAFE cell with respect to its starting position, corresponding to an average between the repetitions of each type of experiment.
Figure 2: Center of mass. Snapshots of experiments (A) GLASS, (B) THICK, (C) THIN and (D) CONTROL (see text for details). Green, blue and red traces correspond to the motion of the center of mass for consecutive $5$-minutes time windows from the start of an experiment (the snapshots were taken at the end of each video). (E-H) Time evolution of the center of mass measured from its starting position along the horizontal direction (i.e., $x-x_{0}$) inside the SAFE (right) cell for each protocol ($x=0$ is taken at the center of the cell). Light-colored shadows indicate standard deviation. Notice that a repellent is added in the DANGER (left) cell at $t=5$ minutes into the repellent reservoir seen as a white spot in (D).
@@
Figure 2C,G also reveals an interesting behavior of the ants in the SAFE arena as they escape from the barrier. Immediately following the repellent injection, the velocity of the CM rapidly spikes, indicating a fast dynamic response that lasts for about one minute. This is followed by a slower process in which the ants form a stable cluster at a certain distance from the barrier, a retreat that is quite distinct from the immediate reaction seen when ants are directly exposed to insect repellent Altshuler et al. (2005). Further details of this behavior will be discussed later.
@@ V Motion analysis before panic
Figure 3: Suppressing boundary effects. (A) Acceleration as a function of position. The inset plot shows acceleration absolute value as a function of the distance from the walls. A characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ is defined as the effective wall thickness. (B) The grey sections have been removed from a sample trajectory, as they are within a distance $\ell_{w}$ from the walls: in order to ensure that statistics is not contaminated by boundary effects, we will only use for future analysis the non-gray sections.
@@
Data consists in approximately $20$ randomly selected ants that were tracked manually during different repetitions of the CONTROL protocol. Trajectories were taken at intervals of $1/3$ of a second, so position samples are spaced at intervals of $\Delta t\approx 0.3\,\mathrm{s}$.
Figure 4: Experimental results before panic. (A) Mean (blue and orange) and fluctuation intensity (dark and light green) of parallel and perpendicular acceleration components for a given speed value. The mean parallel acceleration scatters linearly while perpendicular components remains at zero. Both components of standard deviation are equal and speed-dependent. The fitting line is an exponential of the form of Equation (4b). Although they are plotted on the same axes, note that $\mu(v)$ and $\sigma(v)$ have different units (i.e., $\mathrm{cm\,s}^{-2}$ and $\mathrm{cm\,s}^{-3/2}$, respectively). (B) Velocity autocorrelation function (VACF) up to a lag time of $2\,\mathrm{s}$. The exponential decay is governed by a persistence time $\gamma^{-1}=0.53\,\mathrm{s}$. (C) Average of the local flatness during different time intervals (see Equation (5)). The dashed line indicates the flatness of the Gaussian. In the inset plot, a typical realization of the diffusion coefficient, $D(t)$ (dark red), shows how it changes slowly compared to $v_{x}(t)$ (dark blue). Both quantities are plotted as a function of time over approximately 50 seconds, though neither time nor value ticks are displayed for clarity. (D) Probability density of the diffusion coefficient $D$ (in units of its standard deviation) extracted from velocity time series as indicated in Equation (6). The data was found to fit a Log-normal distribution. For a better visualization of the tail fit, a logarithmic plot is shown as an inset up to $15$ standard deviations.
To eliminate boundary effects, we measured first the extent of wall “forces”. Hence, experimental acceleration components (i.e., $a_{x}$ and $a_{y}$) before repellent injection, were arranged based on increasing values of $x$- and $y$- position components, respectively. Both acceleration time series were binned and a running average was taken (see Figure 3A). The ants slow down as they approach the walls, revealing the influence of the boundaries as if they were confined by a soft, square-box-like potential. Consequently, a plot of acceleration absolute value against the distance from the wall was made, and a characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ resulted from fitting to an exponential function (inset of Figure 3A). Henceforth, the analysis in the rest of this section is conducted with the sections of all trajectories that were consistently at least $\ell_{w}$ units away from every boundary, to discard boundary effects.
@@
we obtain a new coordinate system that changes direction with $\mathbf{v}(t)$, where we projected the instantaneous acceleration. Note that $v\equiv|\mathbf{v}|$.
Figure 4A shows that the acceleration parallel to the velocity at any time $t$ is a random vector whose average is proportional to $-v$, while orthogonal accelerations locally average to zero for all speed values. The standard deviation of both components, however, are not independent of velocity, as demonstrated in the upper part of the same figure, and within their error bars, they exhibit the same non-linear dependence on $v$. Therefore, for the mean of the parallel component, $\langle\mathbf{a}^{(\|)}\rangle$, and the intensity of both the parallel and perpendicular fluctuations, $\langle(\mathbf{a}^{(\|,\bot)}-\langle\mathbf{a}^{(\|,\bot)}\rangle)^{2}\rangle^{1/2}\Delta t^{1/2}$, which we denote as $\mu(v)$ and $\sigma(v)$, respectively, we propose the following relations:
@@
where $\gamma^{-1}$ is a persistence time. The first term, $\Sigma$, accounts for the intensity of acceleration fluctuations at rest, while the second term describes its increase with the velocity, mediated by a characteristic speed $v_{\sigma}$ and a proportionality factor, $\Delta\Sigma$. Note that, in order to obtain $\sigma(v)$, the standard deviation of acceleration time series was scaled by $\Delta t^{1/2}$ to ensure dimensional consistency with a continuous stochastic process, where the white noise has units of $\mathrm{s}^{-1/2}$. We also tested other models for the function $\sigma(v)$, such as general power laws of the form $a+b\,v^{c}$, but none performed better than the exponential model in Equation (4b), as they could, for instance, produce unphysical values for $\sigma(v=0)$. Equation (4b) was fitted and the numerical values of the parameters are reported in Table 1. These results strongly suggest to begin with a Langevin equation such as Equation (1) as a model for ant movement, with a slight modification that allows for $\sigma$ to be expressed as $\sigma(v)$, a function of speed.
Table 1: Values for model parameters fitted to data.
$\ell_{w}$ $a_{0}$ $\gamma^{-1}$ $\Sigma$ $\Delta\Sigma$ $v_{\sigma}$
0.6 1.0 0.53 0.34 1.99 1.99
$s$ $\tau_{f}$ $\tau_{s}$ $L$ $t_{p}$ $\tilde{x}$
1.27 20 3.2 6.0 5.5 3.8
Alongside, the temporal correlation of ant velocities decay exponentially, as shown in a plot of the velocity autocorrelation function (VACF) depicted in Figure 4B. With this, we found a persistence time of $\gamma^{-1}=0.53\,\mathrm{s}$ (also reported in Table 1), which is consistent with the inverse of the slope of parallel acceleration plot shown in Figure 4A. Let’s now look in more detail into the stochastic part of accelerations.
Figure 5A shows the probability density function (PDF) for the zero-mean parallel and perpendicular acceleration components, i.e., $\mathbf{a}^{(\|,\bot)}-\langle\mathbf{a}^{(\|,\bot)}\rangle$, in units of its standard deviation. The non-Gaussian shape highlights the need for a model with non-Gaussian noise. A straightforward assumption is to allow the noise intensity to be stochastic. This leads us to the domain of Superstatistics, as the more basic consideration following a constant noise intensity is that it should be a slowly varying process. Note that in Figure S3, the noise autocorrelation function (NACF) appears as a flat function of time close to zero, so acceleration fluctuations are not correlated at the time scales under study.
A dynamical reason for the non-Gaussian behavior seen in Figure 5A has been proven to be due to Superstatistics whenever a superposition of local Gaussians (2) whose variance fluctuates on a large spatio-temporal scale arise. Therefore, it is needed to check if time series does contain separate time scales. For that, as suggested by Beck et al. (2005), a time average of the local flatness, $\kappa_{i}$, of velocity components, $v_{i}$, was computed for increasing time intervals, $\tau$:
@@
Here, $T$ is the overall measurement time and $\langle v_{i}\rangle$ is the average of the $i$-th velocity component on a time window, $\tau$. The result is shown in Figure 4C, where a dashed line was drawn to indicate the flatness of the Gaussian. The time $\tau_{2}$ at which this line intercepts $\kappa_{i}(\tau)$ corresponds to a time scale $\tau_{2}=3.3\,\mathrm{s}$. Having pointed out that the time scale at which velocity fluctuates is $\tau_{1}\equiv\gamma^{-1}=0.53\,\mathrm{s}$, the ratio $\tau_{2}/\tau_{1}\approx 6$ implies that there exist some intensive parameter, $\beta(t)$, that fluctuates on a temporal scale 6 times larger than the one of the velocity, $\mathbf{v}(t)$.
@@
where $\langle\cdot\rangle_{\tau_{2}}$ means that the average is taken over a time window of the order of $\tau_{2}$. So, we looked for a suitable power $p$ such that $\beta^{p}$ could make the dynamics fall into one of the universality superstatistical classes, e.g., $\chi^{2}$ or Log-normal Beck (2006). We found that, indeed, when $p=-1$, it is the diffusion coefficient, $D\propto\beta^{-1}$, the slowly varying intensive parameter. Furthermore, Figure 4D shows that it follows a Log-normal density that, in its standardized form, looks like:
@@
Here, at any time instant $t$, the diffusion coefficient $D(t)$ and the Gaussian $\xi_{i}(t)$ are used to generate the stochastic variable $\eta(t)=\sqrt{2\gamma^{2}D(t)}\,\xi_{i}(t)$. Then $\eta(t)$ is standardized by dividing by its standard deviation to be finally scaled by $\sigma(v)$. A Runge-Kutta method of strong order $1.5$ within a Milstein scheme to compute Stratonovich-type integrals was used for numerical integration Burrage and Burrage (1996).

### v3

@@ Abstract
Nonequilibrium systems with strong parameter fluctuations are challenging to describe with standard Statistical Mechanics techniques. Superstatistics –that can be seen as statistics of an underlying family of statistical distributions– has emerged as a tool able to describe these complex systems. This has been successfully applied, for example, to hydrodynamic turbulence, internal convection, and even DNA architecture, but not to macroscopic biological systems. Here we document the occurrence of Superstatistics in the animal world, and reveal the emergence of Log-normal Superstatistics in a living system, when ants in a confined space are exposed to a threat. We use a data-driven superstatistical model to explain both normal and “panic” dynamics, identifying non-Gaussian velocity distributions, time scale separation, and the Log-normal statistics of a stochastic diffusion coefficient. We also reveal distinct behavioral regimes in the ants’ collective panic response and how it relates with the individual ant memory and cluster formation. Furthermore, our findings indicate that optical signals or simple antennation are not significant mechanisms for panic transmission. These discoveries provide a foundation for the understanding of the biological origin of Log-normal type diffusion in confined environments.
@@
The authors declare no conflict of interest.
Author contributions: E.A., M.S.T. and O.R. designed the experiments. A.Re., M.C., F.T., E.A. and A.Ri. performed the experiments. A.Re. and M.C. worked on manual tracking. A.Re. analyzed the data and developed the theoretical model. A.Re., M.C. and E.A. wrote the manuscript.
@@ I Significance Statement
We show that ants, when confined and exposed to a threat, display collective movement behaviors that challenge traditional Statistical Mechanics. We provide evidence of Superstatistics in the animal kingdom, highlighting how fluctuations in their dynamics lead to non-Gaussian velocity distributions and Log-normal diffusion patterns. These findings identify distinct behavioral regimes linked to memory and clustering without significant reliance on optical or antennal signaling. This work demonstrates how statistical models can describe the complexity of biological systems, providing insights into the mechanisms driving collective responses in confined environments and offering potential applications in areas like swarm robotics or evacuation modeling.
@@ II Introduction
Nature displays a diversity of collective biological behaviors ranging from the microscopic to the macroscopic scales, and from high confined environments to the open space: the movement of cells during wound healing Vedel et al. (2013); Angelini et al. (2011); Zhang et al. (2010) and the flocking of birds Cavagna et al. (2010); Ballerini et al. (2008); Pearce et al. (2014) illustrate quite well those scenarios.
Here, we focus on the collective motion of confined macroscopic individuals, which is a common scenario in eusocial organisms, such as ants, who often live in high density societies. Leaf-cutter ants like Atta Insularis build complex nests made of networks of tunnels and chambers within which they live, move, take care of eggs, and even harvest fungus. While there is an extensive body of work that qualitatively describes the activity of ants in both open and confined spaces, detailed quantitative models are still needed Hölldobler and Wilson (1990); Gordon (2010); Sumpter (2006).
In terms of quantitative modelling descriptions based on a Langevin equation have been proposed, which captures the finding that animals tend to maintain their direction of motion during some characteristic (persistence) time Méndez et al. (2016); Viswanathan et al. (2011); Gurarie et al. (2017); Santana-Filho et al. (2020):
@@
where $v_{i}$ denotes the $i$-th Cartesian component ($i=x,y$) of the velocity vector, $\gamma$ is a positive parameter that encloses directional persistence, and $\sigma$ is the intensity of a Gaussian white noise, $\xi_{i}$. A more complete overview of the statistical properties of Equation (1) was made in 1931 and it is known as Ornstein-Uhlenbeck process Uhlenbeck and Ornstein (1930), nowadays a reference in ecological models for persistent movement Méndez et al. (2016).
In addition to this, a superposition of different dynamics on different time scales often describes properties of driven nonequilibrium systems Beck (2001), e.g., the fast jiggly dynamics of a Brownian particle subject to slow environmental fluctuations of a given diffusivity Wang et al. (2009); Chechkin et al. (2017). This is known as Superstatistics (which stands for superposition of statistics), and one of its key strengths is the ability to explain the emergence of non-Gaussian densities Beck and Cohen (2003); Metzler (2020).
The theory of Superstatistics is connected to Langevin systems and, within the framework of Statistical Mechanics, can be interpreted by considering a system that is composed of small regions or cells that are temporarily in local equilibrium, showing spatio-temporal fluctuations of some intensive quantity, such as the inverse temperature $\beta=2\gamma/\sigma^{2}$ in Equation (1). It is assumed that the system reaches local equilibrium on a timescale $\tau_{1}$ much smaller than the time $\tau_{2}$ needed for $\beta$ to vary among cells. Consequently, within each region, $\beta$ could be considered approximately constant. Then, during time intervals of the order of $\tau_{2}$ and inside each cell of a system following the dynamics of Equation (1), local Gaussian velocity stationary distributions of the form:
@@
should be computed. Here, $f(\beta)$ is the probability density function of $\beta$, and a relation such as Equation (3) helps us to understand the emergence of non-Gaussian behavior reported in the literature Queirós (2008); Rizzo et al. (2004); Porporato et al. (2006); Briggs and Beck (2007); Chen and Beck (2008).
Through experiments, here we use the behavior of confined ants in normal and panic conditions as a case study for superstatistical analysis. We quantify the motion of ants that are confined into a squared, quasi-2D cell. A number of interesting features arise: from non-Gaussian ant velocity distribution heavily influenced by speed-dependent noise intensity, to the existence of a fluctuating diffusivity, characteristic of non-Gaussian diffusion emergence in single-particle tracking experiments performed in soft and biological matter Wang et al. (2012); Chen and Beck (2008); Devereux et al. (2021). Then, we expose the confined ants to the effect of danger information: ants in a neighbouring cell that are exposed to an unpleasant chemical agent transmit “panic information” to the individuals under study via a specially-designed barrier through which the ants cannot physically cross. Here, we refer to panic as a state of enhanced mobility associated with a heightened state of defensive behavior. A collective output emerges within minutes: in spite of their short individual memory Gordon (2010), the ants form a cluster whose center shifts away from the barrier through which the danger information has emerged. In other words, the individual short-term memory has become a long-term collective memory when danger information is turned on. We propose a superstatistical model able to quantitatively reproduce all the experimentally observed features, both before and after panic. To the best of our knowledge, this is the first documented occurrence of Superstatistics in the animal world and Log-normal Superstatistics in any living system.
@@ III Experimental details
A system consisting in two 3D-printed cells or arenas, each of $6\times 6\times 1\,\mathrm{cm^{3}}$, covered with a glass sheet and separated by a barrier, was designed to investigate the possible transmission of danger information by ants from one cell to another (see Figure 1). Ants are placed in both cells and a video is recorded for $5$ minutes. Then, an insect repellent is added at the center of the left cell and another $10$ minutes of video is recorded. The left and right cells will be referred to as danger and safe, respectively. In a CONTROL experiment the same procedure is followed, except no ants are introduced into the danger arena. We hypothesize that, if ants in the safe arena move away from the barrier after the addition of the repellent a significantly larger distance than they do during a CONTROL experiment, danger information has been transmitted by the “panicked” ants in the danger cell to the ants in the safe cell beyond any “passive” diffusion of the repellent through the barrier.
@@
In an attempt to isolate the possible mechanisms of danger information transmission, various types of barriers were employed (lower panel of Figure 1A). The glass barrier consists of a transparent glass wall that allows visual signals to propagate but filters out all physical interactions. The thick barrier is made in such a way that ants on both sides can make contact with their extremities (e.g., their respective antennas) but cannot antennate the bodies of individuals in the opposite arena. In contrast, the thin barrier allows such contacts (see Figure 1C). We term our protocols GLASS, THICK and THIN, corresponding to each type of barrier.
@@ IV Collective motion: Tracking the Center of Mass
We tracked the center of mass (CM) in both cells for each protocol. Data was processed using yupi, a Python library to handle (tracking, transform, generate, store, analyze, visualize) trajectory data Reyes et al. (2023). As illustrated in Figure 2A-C for typical experiments, the motion of the CM at the danger cell is always near its geometrical center, independently from the protocol used. A similar situation holds for the safe cell, with the conspicuous exception of THIN, where the CM tends to move away from the barrier after the repellent is added into the danger arena. Figure 2E-H shows the time evolution of the $x$-coordinate of the CM at the safe cell with respect to its starting position, corresponding to an average between the repetitions of each type of experiment.
Figure 2: Center of mass. Snapshots of experiments (A) GLASS, (B) THICK, (C) THIN and (D) CONTROL (see text for details). Green, blue and red traces correspond to the motion of the center of mass for consecutive $5$-minutes time windows from the start of an experiment (the snapshots of ants were taken at the end of each video). (E-H) Time evolution of the center of mass measured from its starting position along the horizontal direction inside the safe (right) cell for each protocol ($x=0$ is taken at the center of the cell). Light-colored shadows indicate standard deviation. In all experiments, repellent is added in the danger (left) cell at $t=5$ minutes into the repellent reservoir seen as a white spot in (D).
@@
Figure 2G also reveals an interesting behavior of the ants in the safe arena as they escape from the barrier. Immediately following the repellent injection, the velocity of the CM rapidly spikes, indicating a fast dynamic response. This is followed by a slower process in which the ants form a stable cluster at a certain distance from the barrier, a retreat that is quite distinct from the immediate reaction seen when ants are directly exposed to insect repellent Altshuler et al. (2005). Further details of this behavior will be discussed later.
@@ V Motion analysis before panic
Data consists in 18 randomly selected ants that were tracked manually during different repetitions of the CONTROL protocol. To eliminate boundary effects, we measured first the extent of wall “forces”. Hence, experimental acceleration components (i.e., $a_{x}$ and $a_{y}$) before repellent injection, were arranged based on increasing values of $x$- and $y$- position components, respectively. Both acceleration time series were binned and a running average was taken (see Figure 3A). The ants slow down as they approach the walls, revealing the influence of the boundaries as if they were confined by a soft, square-box-like potential. Consequently, a plot of acceleration absolute value against the distance from the wall was made, and a characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ resulted from fitting to an exponential function (inset of Figure 3A). Henceforth, the analysis in the rest of this section is conducted with the sections of all trajectories that were consistently at least $\ell_{w}$ units away from every boundary, to discard boundary effects.
Figure 3: Boundary effects. (A) Acceleration as a function of position (component wise). The inset plot shows acceleration absolute value as a function of the distance from the walls. A characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ is defined as the effective wall thickness. (B) The grey sections have been removed from a sample trajectory, as they are within a distance $\ell_{w}$ from the walls: in order to ensure that statistics is not contaminated by boundary effects, we will only use for future analysis the non-gray sections.
@@
we obtain a new coordinate system that changes direction with $\mathbf{v}(t)$, onto which we project the instantaneous acceleration. Note that $v\equiv|\mathbf{v}|$.
We computed, as a function of speed values, the mean of the parallel acceleration component, $\mu(v)\equiv\langle\mathbf{a}^{(\|)}\rangle$, and the intensity of both the parallel and perpendicular fluctuations, $\sigma(v)\equiv\langle(\mathbf{a}^{(\|,\bot)}-\langle\mathbf{a}^{(\|,\bot)}\rangle)^{2}\rangle^{1/2}\Delta t^{1/2}$, and plotted them in Figure 4A. To obtain $\sigma(v)$, the standard deviation of acceleration time series was scaled by $\Delta t^{1/2}$ to ensure dimensional consistency with a continuous stochastic process, where the white noise has units of $\mathrm{s}^{-1/2}$.
Figure 4: Experimental results before panic. (A) Mean (blue and orange) and fluctuation intensity (dark and light green) of parallel and perpendicular acceleration components, respectively, for a given speed value. The mean parallel acceleration scatters linearly while perpendicular components remains at zero. Both components of standard deviation are equal and speed-dependent. The fitting line is an exponential of the form of Equation (4b). Although they are plotted on the same axes, note that $\mu(v)$ and $\sigma(v)$ have different units (i.e., $\mathrm{cm\,s}^{-2}$ and $\mathrm{cm\,s}^{-3/2}$, respectively). (B) Velocity autocorrelation function (VACF) up to a lag time of $2\,\mathrm{s}$. The exponential decay is governed by a persistence time $\gamma^{-1}=0.53\,\mathrm{s}$. (C) Average of the local flatness during different time intervals (see Equation (5)). The dashed line indicates the flatness of the Gaussian. In the inset plot, a typical realization of the diffusion coefficient, $D(t)$ (dark red), shows how it changes slowly compared to $v_{x}(t)$ (dark blue). Both quantities are plotted as a function of time over approximately 50 seconds, though neither time nor value ticks are displayed for clarity. (D) Probability density of the diffusion coefficient $D$ (in units of its standard deviation) extracted from velocity time series as indicated in Equation (6). The data was found to fit a Log-normal distribution. For a better visualization of the tail fit, a logarithmic plot is shown as an inset up to $15$ standard deviations.
Figure 4A illustrates that the acceleration component parallel to the velocity at any time $t$ is a random vector whose average is proportional to $-v$, while the orthogonal accelerations locally average to zero for all speed values. However, the intensity of both components is not independent of velocity. As demonstrated in the upper part of the same figure, and within their error bars, the intensities exhibit the same non-linear dependence on $v$.
Based on these observations, we propose the following relations:
@@
where $\gamma^{-1}$ in (4a) is a persistence time, the first term in (4b), $\Sigma$, accounts for the intensity of acceleration fluctuations at rest, and the second term describes its increase with the velocity, mediated by a characteristic speed $v_{\sigma}$ and a proportionality factor, $\Delta\Sigma$. We ended with an exponential function because of its simplicity, but we also tested other models for the function $\sigma(v)$, such as general power laws of the form $a+b\,v^{c}$, but none performed better than the exponential in Equation (4b), as they could, for instance, produce unphysical values for $\sigma(v=0)$ (i.e., $\sigma(0)<0$). Equation (4b) was fitted and the numerical values of the constants $\gamma$, $\Sigma$ and $\Delta\Sigma$ are reported in Table 1. These results strongly suggest to begin with a Langevin equation such as Equation (1) as a model for ant movement, with a slight modification that allows for $\sigma$ to be expressed as $\sigma(v)$, a function of speed.
Furthermore, the temporal correlation of ant velocities decay exponentially, as shown in a plot of the velocity autocorrelation function depicted in Figure 4B. With this, we found a persistence time of $\gamma^{-1}=0.53\,\mathrm{s}$ (also reported in Table 1), which is consistent with the inverse of the slope of parallel acceleration plot shown in Figure 4A. Let’s now look in more detail into the stochastic part of accelerations.
Figure 5A shows the probability density function (PDF) for the zero-mean parallel and perpendicular acceleration components, i.e., $\mathbf{a}^{(\|,\bot)}-\langle\mathbf{a}^{(\|,\bot)}\rangle$, in units of its standard deviation. The averages were computed by binning over speed intervals, following the same approach used in Figure 4A. The non-Gaussian shape highlights the need for a model with non-Gaussian noise. A straightforward assumption is to allow the noise intensity to be stochastic. This leads us to the domain of Superstatistics, as the more basic consideration following a constant noise intensity is that it should be a slowly varying process. Note that in Figure S3, the noise autocorrelation function appears as a flat function of time close to zero, so acceleration fluctuations are not correlated at the time scales under study.
A dynamical reason for the non-Gaussian behavior seen in Figure 5A has been proven to be due to Superstatistics whenever a superposition of local Gaussians (2) whose variance fluctuates on a large spatio-temporal scale arise. Therefore, it is needed to check if time series does contain separate time scales. For that, as suggested by Beck et al. (2005), a time average of the local flatness, $\kappa_{i}$, of velocity components, $v_{i}$, was computed for increasing time intervals, $t$:
@@
Here, $T$ is the overall measurement time and $\langle v_{i}\rangle$ is the average of the $i$-th velocity component on a time window, $t$. The goal is to determine the time scale, $\tau_{2}$, at which the velocity time series is sampled from a Gaussian density such as Equation (2). The computed $\kappa_{i}(t)$ is shown in Figure 4C, where a dashed line was drawn to indicate the flatness of the Gaussian. The time $\tau_{2}$ at which this line intercepts $\kappa_{i}(t)$ corresponds to a time scale $\tau_{2}=3.3\,\mathrm{s}$. Having pointed out that the time scale at which velocity fluctuates is $\tau_{1}\equiv\gamma^{-1}=0.53\,\mathrm{s}$, the ratio $\tau_{2}/\tau_{1}\approx 6$ implies that there exist some intensive parameter, $\beta(t)$, that fluctuates on a temporal scale 6 times larger than the one of the velocity, $\mathbf{v}(t)$.
@@
where $\langle\cdot\rangle_{\tau_{2}}$ means that the average is taken over a time window of $\tau_{2}$ units. So, we looked for a suitable power $p$ such that $\beta^{p}$ could make the dynamics fall into one of the universality superstatistical classes, e.g., $\chi^{2}$ or Log-normal Beck (2006). We found that, indeed, when $p=-1$, it is the diffusion coefficient, $D\propto\beta^{-1}$, the slowly varying intensive parameter. Furthermore, Figure 4D shows that the diffusion coefficient follows a Log-normal density that, in its standardized form, looks like:
@@
We refer to Equation (8) as the SSModel, where at any time instant $t$, the diffusion coefficient $D(t)$ and the Gaussian $\xi_{i}(t)$ are used to generate the stochastic variable $\eta(t)=\sqrt{2\gamma^{2}D(t)}\,\xi_{i}(t)$. Then $\eta(t)$ is standardized by dividing by its standard deviation to be finally scaled by $\sigma(v)$. A Runge-Kutta method of strong order $1.5$ within a Milstein scheme to compute Stratonovich-type integrals was used for numerical integration Burrage and Burrage (1996).